# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v2

@@ 3.1.3 Image generation
As described in Section 2.3, recent advances in AI technologies for image generation are based on Diffusion Models (DMs). Well-known and highly competitive text-to-image models include Stable Diffusion [^21] https://stability.ai/stable-image, Midjourney [^22] https://www.midjourney.com/home, DALL·E [^23] https://openai.com/dall-e-3, and Ideogram [^24] https://ideogram.ai/. Released in March 2024, the latest version, Stable Diffusion 3 (SD3) [52], reports outperforming state-of-the-art text-to-image generation systems such as DALL·E 3 (released August 2023), Midjourney v6 (released December 2023), and Ideogram v1 (released February 2024) in terms of typography and prompt adherence, based on human preference evaluations. These open-source tools are built with the Multimodal Diffusion Transformer (MM-DiT) architecture, which integrates attention from both text and image. Examples of text-to-image generation are shown in Fig. 3 (a) comparing the performance of four models, i.e. Ideogram v1, DALL·E 3, Photoshop 2024, and sdxy-turbo by Nvidia. It is obvious that hands are the most difficult to generate, e.g., one hand has six fingers.
@@
Figure 3: Text-to-image generation (generated on 27 November 2024). (a) text-to-image generation by Ideogram v1, DALL·E 3, Photoshop 2024, and sdxy-turbo by Nvidia. (b) The top-row images were generated by DALL·E in ChatGPT 4. The bottom-row images are generated by LLM-grounded Diffusion [56].
@@
DALL·E 3, available on ChatGPT 4, also provides an inpainting tool, allowing the user to manually select the area to edit. However, as of April 2024, the performance is still limited. As illustrated in Fig. 3 (b), the selected area is the white car, and with the follow-up request to change the white car to the red car, DALL·E 3 generates correctly. However, if asked to replace it with a bicycle, it does not work. Recently, LLM-grounded Diffusion [56] enables instruction-based scene specification with multiple rounds of user requests without a need of manual selection on the image. This is achieved by generating layout-grounded images first using stable diffusion, then masking the latent variables as priors for the next round of generation [^25] Images in Fig. 3 (b) were generated using their demo: https://huggingface.co/spaces/longlian/llm-grounded-diffusion..
@@
Similar to DALL·E 3, Photoshop features a Generative Fill tool [^26] https://www.adobe.com/th_en/products/photoshop/generative-fill.html designed to generate new images or assist with photo editing. It accepts a text prompt and provides several generation choices. After defining the editing area, users can remove and add new objects (more inpainting tasks are discussed in Section 3.3.5), transfer to new styles, and expand content within images. Recently, InstructPix2Pix [53] proposed a conditional diffusion model to generate image editing examples without predefined editing areas. It combines GPT-3 and Stable Diffusion. This clearly demonstrates that the model effectively captures and matches the semantic meaning of the content in both text and image. Sometimes, style and context are not easy to describe in words. Textual Inversion [54] personalizes large pre-trained text-to-image diffusion models based on specific objects and styles, using 3-5 images of a user-provided concept. Recently, ByteDance announces Hyper-SD [57] proposed trajectory segmented consistency distillation that provides real-time high-resolution image generation from drawing with a control text prompt.
@@ 3.1.4 Video generation and animation
Despite the success of text-to-image generation, text-to-video generation has just started to grow rapidly in 2024, largely due to its computational expense and content complexity. However, several major companies and private platforms have released their progress in this area, including Gemini 1.5 by Google, Make-A-Video by Meta, and Sora by OpenAI. Make-A-Video [69], through a spatiotemporally factorized diffusion model, leverages joint text-image priors and super-resolution in space and time. Some results however contain flickering artifacts [^27] https://makeavideo.studio/. Gen-2 by Runway [^28] https://research.runwayml.com/gen2 offering both text- and image-to-video can generate a smooth 4-sec video. In April 2024, Adobe Premier Pro has announced their integration of generative AI tools for video extension with third-party models by OpenAI, Runway and Pika Labs [^29] https://www.adobe.com/products/premiere/ai-video-editing.html. Its new update also includes contextual-selection tool, inpainting for object removal, and object addition to the defined areas in the videos with a text prompt.
@@
Text-to-video technologies, combined with AI voice, have been tested not only by artists or producers but also by a wider audience. Results from these tests, such as automatically turning scripts into movie trailers and music videos, have been widely shared on public online platforms [^30] https://twitter.com/minchoi/status/1775907105813217398. However, scene composition and transitions still require further editing to align with producers’ needs [^31] See an example by Curious Refuge at https://www.youtube.com/watch?v=fJQbP34GoHQ. In April 2024, Microsoft introduced VASA-1 [64], which turns a single static image and a speech audio clip into a video clip of realistic talking faces mimicking human facial expressions and head movements, shown in Fig. 4 (right). The overall quality of the generated videos is better than VLOGGER by Google [65], which is based on similar technology – diffusion models. However, VLOGGER also offers movement of the upper body and hand gestures. Recently, ByteDance introduced an audio-driven interactive head generation [68] that offers listening and speaking states during multi-turn conversations. This framework is based on a conditional diffusion transformer . The main technologies of text-to-video and image-to-video tasks are based on DMs with a combination of 3D convolutions (or separately spatial and temporal convolutions), and spatial and temporal attention modules [71]. Tune-A-Video [72] modifies the style of an input video using a text prompt. The method leverages pretrained text-to-image models and introduces attention tuning to ensure temporal consistency. Early video generation methods often exhibit flickering, as observed in the CVPR2023 competition on text-guided video editing, where all results suffered from temporal inconsistency. Dreamix [70] videos do not have this issue, but they are very blurry. As an example of a transformer-based approach, CogVideo [58] employs VQ-VAE to convert input frames to tokens, which are then fused with text tokens to produce a new video. Phenaki [59] exploits transformers to generate variable length videos, but the quality is lower than those based on DMs. Evaluations of these methods can be found in [62]. The recent work applies spatiotemporal layers to model temporal dynamics [66]. The transformer blocks have been redesigned for latent video diffusion modeling with window-restricted spatial and spatiotemporal attention. By the end of 2024, there will be many more free and commercial tools for video generation. These include Veo 2 by Google DeepMind, Kling AI [^32] https://www.klingai.com/, Pika 2.0333333https://pikartai.com/pika-2-0/, Hailuo AI [^34] https://hailuoai.video/, etc. Though not perfect, the generated videos are close to reality (visit their websites for showcase examples).
@@
Viggle [^36] https://viggle.ai/ claims to be the first video-3D foundation model with an actual understanding of physics. It combines a character and a text prompt about motion to generate character animation. Available AI tools on the market include DeepMotion [^37] https://www.deepmotion.com/ that offers text-to-3D post animation and video-to-3D post animation, shown in Fig. 4 (left). The later function can track multiple people from real video and generates replicated characters with the same motions.
@@ 3.1.5 Augmented, virtual and mixed reality, and 3D content
While the benefits of LLMs in Augmented Reality (AR) directly target educational purposes, enhance cognitive support, and facilitate communication [242], mixed reality (MR) has once again become exciting since the release of the Apple Vision Pro in February 2024. It has demonstrated the potential of MR experiences by merging real-world environments with computer-generated ones. Thanks to the rapid growth of AI-based 3D representation (see Section 3.5), the generation of AR/VR/MR content has advanced significantly. Real-time rendering with immersive interaction has improved, and real scenes can now be generated avoiding uncanny valley effects. There has also been an attempt to use autoregressive and generative models to estimate lighting, achieving a visually coherent environment between virtual and physical spaces in AR [79].
Similar to other content generations, LLMs have been influenced on immersive technologies, including text-to-3D and image-to-3D. Exciting examples include Holodeck [73], which automatically generates 3D embodied environments via text prompt conversation of a large language model (GPT-4). 3D objects are gathered from Objaverse [243], a dataset with 800K+ annotated 3D objects. RealFusion [75], a single image to 3D object generator, merges 2D diffusion models with NeRF, improving Instant-NGP [190], which provide API for VR controls. NeuralLift-360 [74] also uses diffusion models to generate prior for novel view synthesis. Magic123 [76] is the latest image-to-3D tool that use 2D and 3D priors simultaneously to produce high-quality high-resolution 3D geometry and textures. DreamGaussian [77] offers text-to-3D and image-to-3D by adapting 3D Gaussian splatting (more in Section 3.5.3) into generative settings using diffusion prior. This generates photo-realistic 3D assets with explicit mesh and texture maps within only 2 minutes. DreamGaussian4D [78] employs image-to-video diffusion and a 4D Gaussian Splatting representation to generate an image-to-4D model. The results are not very sharp, but they can be further edited with Blender. In July 2024, Shutterstock launched its Generative 3D service in commercial beta, powered by NVIDIA Edify, a multimodal generative AI architecture. This service enables creators to rapidly prototype 3D assets and generate 360-degree HDRi backgrounds to light scenes using text or image prompts. In conjunction with OpenUSD, the created scenes can be rendered into 2D images and used as input for AI-powered image generators, allowing for the production of precise, brand-accurate visuals.
@@ 3.2.1 Text categorization
Applications of text categorization include detecting spam emails, automating customer support, monitoring social media for harmful content, etc. At its core, text categorization involves assigning predefined labels to text documents, which can be anything from a tweet to a lengthy article. LLMs are particularly well-suited for this task due to their ability to comprehend complex and nuanced language. One of the main advantages of using LLMs in text categorization is their transfer learning capability. Models can be pre-trained on a large amount of text and then fine-tuned on a smaller, task-specific dataset, with or without further post-processing technique. For example, CARP [80] applies kNN to integrate diagnostic reasoning process for final decision. ChatGraph, proposed by Shi et al. [81], utilizes ChatGPT to refine text documents. It uses a knowledge graph, extracted using another specific defined prompt, and finally, a linear model is trained on the text graph for classification. Multiple learners are also used to enhance the performances [82].
@@ 3.2.2 Advertisements and film analysis
Current advances in generative AI would greatly benefit sentiment analysis, also known as opinion mining, where opinions are gather from social media, articles, customer feedback, and corporate communication and are analysed to understand emotion of the owners. This is a potential tool for filmmakers and studios, enabling the creation of effective and targeted marketing campaigns. By analyzing viewer emotions and opinions, AI can provide valuable insights into audience preferences, aiding in the optimization of film marketing strategies. Sentiment analysis with modern generative AI produce more accurate results. Technically, LLMs learn complex patterns and relationships in text data for sentiment classification [83, 84]. SiEBERT [85] provides pre-trained model with open-source scripts to be fine-tuned to further improve accuracy for novel applications.
@@ 3.2.3 Content retrieval and recommendation services
Generative retrieval (GR) was pioneered by Metzler et al. [86]. Unlike traditional retrieval, which adheres to the “index-retrieve-then-rank” paradigm, the GR paradigm employs a single model to obtain results from query input. The model generally involve transformers, generating output token-by-token. More recent work in [91] introduces learning-to-rank training to enhance the performance system up typ 30%. GR has several advantages including substituting the bulky external index with an internal index (i.e., model parameters), significantly reducing memory usage, and enabling optimization during end-to-end model training towards a universal objective for information retrieval tasks. Conversational question answering techniques have been integrated to enhance the document retrieval [90].
When retrieving visual content, recent work exploits generative models to enhance content-based model search [88]. These models decode the text, image, or video query into samples of possible outputs, which are then used to learn statistics for better matching between the query and output candidates. DMs are also employed for visual retrieval tasks, where they learn joint data distributions between text queries and video candidates [92].
While the retrieval task involves users directly defining a specific query input, recommendation services operate by retrieving content based on previous usage patterns. Essentially, a recommendation engine is a system that suggests products, services, or information to users through data analysis. Research in [245] has reported a positive association between buyers’ attitudes toward AI and their behavioral intention to accept AI-based recommendations, with potential for further growth. Notable examples include the recommendation framework developed by Google [89], which utilizes GR. This framework assigns Semantic IDs to each item and trains a retrieval model to predict the Semantic ID of an item that a given user may engage with.
@@ 3.2.4 Intelligent assistants
Current LLMs obviously enhance the performance of intelligent assistants, designed to understand complex inquiries and generate more natural conversational responses, such as Sasha [93]. Generative AI can also be used to enhance the performance of human customer support agents, aiding in search and summarization, as discussed in the previous section. Brynjolfsson et al. [246] examined the implementation of a generative AI tool designed to offer conversational guidance to customer support agents. Their research revealed that AI assistance significantly enhances problem resolution and customer satisfaction. Furthermore, they observed that AI recommendations prompt low-skill workers to adopt communication styles akin to those of high-skill workers. AI-based intelligent assistants may currently be more focused on educational purposes, but they can clearly help artists write more efficiently [247] or assist in customizing personal requirements [248].

### v3

@@ 3.1.3 Image generation
As described in Section 2.3, recent advances in AI technologies for image generation are based on Diffusion Models (DMs). Well-known and highly competitive text-to-image models include Stable Diffusion [^25] https://stability.ai/stable-image, Midjourney [^26] https://www.midjourney.com/home, DALL·E [^27] https://openai.com/dall-e-3, and Ideogram [^28] https://ideogram.ai/. Released in March 2024, the latest version of Stable Diffusion (SD3), has been reported to outperform state-of-the-art text-to-image generation systems such as DALL·E 3 (released August 2023) [51], Midjourney v6 (released December 2023), and Ideogram v1 (released February 2024) in terms of typography and prompt adherence, based on human preference evaluations. These open-source tools are built on a Multimodal Diffusion Transformer (MM-DiT) architecture, which integrates attention from both text and images. LLM4GEN [58] fuses features from LLM and CLIP models to enhance the semantic understanding in text-to-image diffusion models, enabling them to better handle complex and dense prompts involving multiple objects. Examples of text-to-image generation are shown in Fig. 3 (a) comparing the performance of four models, i.e. Ideogram v1, DALL·E 3, Photoshop 2025, and sdxy-turbo by Nvidia. It is clear that hands are one of the most difficult features to generate, e.g., one hand has six fingers.
@@
Figure 3: Text-to-image generation (generated on 27 November 2024). (a) text-to-image generation by Ideogram v1, DALL·E 3, Photoshop 2025, and sdxy-turbo by Nvidia. (b) The top-row images were generated by DALL·E in ChatGPT 4. The bottom-row images are generated by LLM-grounded Diffusion [55].
@@
DALL·E 3, available on ChatGPT 4, also provides an inpainting tool, allowing the user to manually select the area to edit. However, as of April 2024, its performance is still limited. As illustrated in Fig. 3 (b), the selected area is the white car, and with the follow-up request to change the white car to the red car, DALL·E 3 generates correctly. However, if asked to replace it with a bicycle, it does not work. LLM-grounded Diffusion [55] was the first to introduce a framework that allows multiple rounds of user requests without the need for manual selection on the image. This is achieved by generating layout-grounded images, first using stable diffusion and then masking the latent variables as priors for the next round of generation [^29] Images in Fig. 3 (b) were generated using their demo: https://huggingface.co/spaces/longlian/llm-grounded-diffusion.. Since then, text-driven image editing has seen significant improvements in quality, with most recent approaches adopting Diffusion Transformer architectures [57, 267].
@@
Similar to DALL·E 3, Photoshop features a Generative Fill tool [^30] https://www.adobe.com/th_en/products/photoshop/generative-fill.html designed to generate new images or assist with photo editing. It accepts a text prompt and provides several generation choices. After defining the editing area, users can remove and add new objects (more inpainting tasks are discussed in Section 3.3.5), transfer to new styles, and expand content within images. Recently, Brooks et al. introduced InstructPix2Pix [52], a conditional diffusion model that generates image editing examples without predefined editing areas. By combining GPT-3 and Stable Diffusion, the model effectively captures and matches the semantic meaning of the content in both text and image. Sometimes, style and context are not easy to describe in words. Textual Inversion [53] personalizes large pre-trained text-to-image diffusion models based on specific objects and styles, using 3-5 images of a user-provided concept. ByteDance announced Hyper-SD [56] which proposed trajectory segmented consistency distillation and provides real-time high-resolution image generation from drawing with a control text prompt.
@@ 3.1.4 Video generation and animation
Despite the success of text-to-image generation, text-to-video generation has not advanced at the same pace, only starting to grow more rapidly in 2024, largely due to its computational expense and content complexity. Several major companies and private platforms have however now released offerings, including Gemini 1.5 by Google, Make-A-Video by Meta, and Sora by OpenAI. Make-A-Video [70], through a spatiotemporally factorized diffusion model, leverages joint text-image priors and super-resolution in space and time. Some results however contain flickering artifacts [^31] https://makeavideo.studio/. Gen-2 by Runway [^32] https://research.runwayml.com/gen2 offers both text- and image-to-video and can generate a smooth 4-sec video. In April 2024, Adobe Premier Pro announced their integration of generative AI tools for video extension with third-party models by OpenAI, Runway and Pika Labs [^33] https://www.adobe.com/products/premiere/ai-video-editing.html. This new update also includes a contextual-selection tool, inpainting for object removal, and object addition to the defined areas in the videos with a text prompt.
@@
Text-to-video technologies, combined with AI voice, have been tested not only by artists or producers but also by a wider audience. Results from these tests, such as automatically turning scripts into movie trailers and music videos, have been widely shared on public online platforms [^34] https://twitter.com/minchoi/status/1775907105813217398. However, scene composition and transitions still require further editing to align with producers’ needs [^35] See an example by Curious Refuge at https://www.youtube.com/watch?v=fJQbP34GoHQ. In April 2024, Microsoft introduced VASA-1 [65], which turns a single static image and a speech audio clip into a video clip of realistic talking faces mimicking human facial expressions and head movements, as shown in Fig. 4 (right). The overall quality of the generated videos is better than VLOGGER by Google [66], which is based on similar technology – diffusion models. However, VLOGGER also offers movement of the upper body and hand gestures. Recently, ByteDance introduced an audio-driven interactive head generation [69] that offers listening and speaking states during multi-turn conversations. This framework is based on a conditional diffusion transformer . The main technologies underpinning text-to-video and image-to-video tasks are based on DMs with a combination of 3D convolutions (or separately spatial and temporal convolutions), and spatial and temporal attention modules [72]. Tune-A-Video [73] modifies the style of an input video using a text prompt. The method leverages pretrained text-to-image models and introduces attention tuning to ensure temporal consistency. Early video generation methods often exhibit flickering, as observed in the CVPR2023 competition on text-guided video editing, where all results suffered from temporal inconsistency. Dreamix [71] videos do not have this issue, but they are very blurry. As an example of a transformer-based approach, CogVideo [59] employs VQ-VAE to convert input frames to tokens, which are then fused with text tokens to produce a new video. Phenaki [60] exploits transformers to generate variable length videos, but the quality is lower than those based on DMs. Evaluations of these methods can be found in [63]. More recent work has applied spatiotemporal layers to model temporal dynamics [67]. The transformer blocks have been redesigned for latent video diffusion modeling with window-restricted spatial and spatiotemporal attention. LaVie [74] demonstrates that simple temporal self-attention mechanisms, when combined with rotary positional encoding, are sufficient to capture the temporal correlations inherent in video data. The image-to-video generation process is analogous to text-to-video methods, but it conditions diffusion models on images rather than text. Some approaches further enhance generation by incorporating both textual descriptions (to guide motion) and images (to define objects and scenes) as inputs [75]. Many more free and commercial tools for video generation are now emerging. These include Veo 3 by Google DeepMind [^36] https://deepmind.google/models/veo/, Kling AI [^37] https://www.klingai.com/, Pika 2.2383838https://pikartai.com/pika-2-2/, Hailuo AI [^39] https://hailuoai.video/, etc. Though not perfect, the generated videos are close to reality (visit their websites for showcase examples).
@@
Viggle [^41] https://viggle.ai/ claims to be the first video-3D foundation model embodying an actual understanding of physics. It combines a character and a text prompt about motion to generate character animation. Available AI tools for 3D on the market include DeepMotion [^42] https://www.deepmotion.com/ that offers text-to-3D post animation and video-to-3D post animation, shown in Fig. 4 (left). The later function can track multiple people from real video and generates replicated characters with the same motions.
@@ 3.1.5 Augmented, virtual and mixed reality, and 3D content
While the benefits of LLMs in Augmented Reality (AR) directly target educational purposes, enhance cognitive support, and facilitate communication [268], mixed reality (MR) has once again become exciting since the release of the Apple Vision Pro in February 2024. This demonstrated the potential of MR experiences by merging real-world environments with computer-generated ones. Thanks to the rapid growth of AI-based 3D representation (see Section 3.5), the generation of AR/VR/MR content has advanced significantly. Real-time rendering with immersive interaction has improved, and real scenes can now be generated avoiding uncanny valley effects. There has also been an attempt to use autoregressive and generative models to estimate lighting, achieving a visually coherent environment between virtual and physical spaces in AR [82].
Similar to other content generation tools, LLMs have been influenced on immersive technologies, including text-to-3D and image-to-3D. Exciting examples include Holodeck [76], which automatically generates 3D embodied environments via text-prompt interactions with a large language model (GPT-4). 3D objects are gathered from Objaverse [269], a dataset with 800K+ annotated 3D objects. RealFusion [78], a single image to 3D object generator, merges 2D diffusion models with NeRF, improving Instant-NGP [209], which provides an API for VR controls. NeuralLift-360 [77] also uses diffusion models to generate priors for novel view synthesis. Magic123 [79] is the latest image-to-3D tool that uses 2D and 3D priors simultaneously to produce high-quality high-resolution 3D geometry and textures. DreamGaussian [80] offers text-to-3D and image-to-3D by adapting 3D Gaussian splatting (more in Section 3.5.3) into generative settings using a diffusion prior. This generates photo-realistic 3D assets with explicit mesh and texture maps within only 2 minutes. DreamGaussian4D [81] employs image-to-video diffusion and a 4D Gaussian Splatting representation to generate an image-to-4D model. The results are not very sharp, but they can be further edited with Blender.
In July 2024, Shutterstock launched its Generative 3D service in commercial beta, powered by NVIDIA Edify, a multimodal generative AI architecture. This service enables creators to rapidly prototype 3D assets and generate 360-degree HDRi backgrounds to light scenes using text or image prompts. In conjunction with OpenUSD, the created scenes can be rendered into 2D images and used as input for AI-powered image generators, allowing for the production of precise, brand-accurate visuals.
@@ 3.2.1 Text categorization
Applications of text categorization include detecting spam emails, automating customer support, monitoring social media for harmful content, etc. At its core, text categorization involves assigning predefined labels to text documents, which can be anything from a tweet to a lengthy article. LLMs are particularly well-suited for this task due to their ability to comprehend complex and nuanced language. One of the main advantages of using LLMs in text categorization is their transfer learning capability. Models can be pre-trained on a large amount of text and then fine-tuned on a smaller, task-specific dataset, with or without further post-processing technique. For example, CARP [83] applies kNN to integrate diagnostic reasoning process for final decision. ChatGraph, proposed by Shi et al. [84], utilizes ChatGPT to refine text documents. It uses a knowledge graph, extracted using another specific defined prompt, and finally, a linear model is trained on the text graph for classification. Multiple learners are also used to enhance the performances [85, 86].
@@ 3.2.2 Advertisements and film analysis
Current advances in generative AI would greatly benefit sentiment analysis, also known as opinion mining, where opinions are gather from social media, articles, customer feedback, and corporate communication and are analysed to understand emotion of the owners. This is a potential tool for filmmakers and studios, enabling the creation of effective and targeted marketing campaigns. By analyzing viewer emotions and opinions, AI can provide valuable insights into audience preferences, aiding in the optimization of film marketing strategies. Sentiment analysis with modern generative AI produce more accurate results. Technically, LLMs learn complex patterns and relationships in text data for sentiment classification [87, 88]. SiEBERT [89] provides pre-trained model with open-source scripts to be fine-tuned to further improve accuracy for novel applications. Cinema Multiverse Lounge [271], a multi-agent conversational system, allows users to interact with LLM-driven agents, each embodying a distinct film-related target users.
@@ 3.2.3 Content retrieval and recommendation services
Generative retrieval (GR) was pioneered by Metzler et al. [90]. Unlike traditional retrieval, which adheres to the “index-retrieve-then-rank” paradigm, the GR paradigm employs a single model to obtain results from query input. The model generally involve deep-learning based transformers, generating output token-by-token. More recent work in [95] introduces learning-to-rank training to enhance the performance system up to 30%. GR has several advantages including substituting the bulky external index with an internal index (i.e., model parameters), significantly reducing memory usage, and enabling optimization during end-to-end model training towards a universal objective for information retrieval tasks. Conversational question answering techniques have been integrated to enhance the document retrieval [94].
When retrieving visual content, recent work exploits generative models to enhance content-based model search [92]. These models decode the text, image, or video query into samples of possible outputs, which are then used to learn statistics for better matching between the query and output candidates. DMs are also employed for visual retrieval tasks, where they learn joint data distributions between text queries and video candidates [96]. A comprehensive survey on Generative Information Retrieval is available in [272].
While the retrieval task involves users directly defining a specific query input, recommendation services operate by retrieving content based on previous usage patterns. Essentially, a recommendation engine is a system that suggests products, services, or information to users through data analysis. Research in [273] has reported a positive association between buyers’ attitudes toward AI and their behavioral intention to accept AI-based recommendations, with potential for further growth. Notable examples include the recommendation framework developed by Google [93], which utilizes GR. This framework assigns Semantic IDs to each item and trains a retrieval model to predict the Semantic ID of an item that a given user may engage with. A report by Aggarwal et al. [274] states that the recommendation accuracy of recommendation services has increased from 45.0% to 91.5% with the integration of generative AI.
@@ 3.2.4 Intelligent assistants
Current LLMs obviously enhance the performance of intelligent assistants, designed to understand complex inquiries and generate more natural conversational responses, such as Sasha [97]. Generative AI can also be used to enhance the performance of human customer support agents, aiding in search and summarization, as discussed in the previous section. Brynjolfsson et al. [275] examined the implementation of a generative AI tool designed to offer conversational guidance to customer support agents. Their research revealed that AI assistance significantly enhances problem resolution and customer satisfaction. Furthermore, they observed that AI recommendations prompt low-skill workers to adopt communication styles akin to those of high-skill workers. AI-based intelligent assistants may currently be more focused on educational purposes, but they can clearly help artists write more efficiently [276] or assist in customizing personal requirements [277]. The performance of personalized assistants can be enhanced with domain-specific knowledge to provide more in-depth responses to users [278].